OpenStudy (anonymous):

Factor completely: c2 - 10c - 24

OpenStudy (anonymous):

sorry, c^2 - 10c - 24

OpenStudy (anonymous):

\[(c-12)(c+2)\]

OpenStudy (saifoo.khan):

right.

OpenStudy (anonymous):

how did you get that? I still have two more to do and I want to try to do them

OpenStudy (anonymous):

could you walk me through it?

OpenStudy (anonymous):

ok so you need two numbers A and B such at AB=the last term in your quadratic and A+B= the middle term in your quadratic given your c^2 has no coefficants

OpenStudy (anonymous):

take your last term break it down into its factoring groups like 24 was 1*24 or 2*12 or 3*8 or 4*6 so i know i can make 10 with two of those (2*12 or 6*4) and since i got a negative in the last term i can rule out 6*4 cause -6+-4 =-10 but -6*-4=24 not -24 so 12*2 is the last option and you arrange the the numbers with the - sign and the postive sign to make that middle term work

OpenStudy (anonymous):

To try to make it more clear: You are looking for some kind of factored form like: (c + a)(c + b) Well if we multiply those two out (using foil or regular distribution) you get: c^2 + ac + bc + ab Then we combine those two middle terms: c^2 + (a+b)c + (a*b) Now compare that to the expression you started with: c^2 + (a+b)c + (a*b) = c^2 -10c - 24 And we can see that a+b = -10 And that a*b = -24 Once we have those two equations we can solve for a and b, and plug them back into our original: (c+a)(c+b) will be the factored form.

OpenStudy (anonymous):

ty